# Next-to-leading order analysis of $J/\psi+\gamma$ production in photon-photon collisions at CEPC

### v1

@@ IV Summary
Within the NRQCD factorization at NLO QCD precision, we systematically investigate $J/\psi+\gamma$ production in $\gamma\gamma$ collisions using three representative long distance matrix element sets that consistently describe $J/\psi$ hadroproduction data. Our results demonstrate absolute dominance of the direct process, exceeding single- and double-resolved contributions by over two orders of magnitude; at CEPC’s design luminosity, this yields $\mathcal{O}(10^{3})$ reconstructed $J/\psi$ events annually. NLO corrections reduce the CS cross section by $\sim$50% while CO states show moderate contributions. For polarization, the CS model predicts pronounced transverse polarization, whereas NRQCD predictions exhibit strong LDME dependence: two sets yield transverse polarization consistent with CSM, while the third generates near-unpolarized to moderately longitudinal behavior – a striking divergence. Furthermore, the polarization predictions depend critically on the ratio $\frac{\langle\mathcal{O}^{J/\psi}(^{3}P_{J}^{[8]})\rangle}{\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[1]})\rangle}$, but not on the specific values of $\langle\mathcal{O}^{J/\psi}(^{1}S_{0}^{[8]})\rangle$ and $\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[8]})\rangle$. Thus, $\gamma\gamma\to J/\psi+\gamma+X$ provides a critical probe for testing LDME universality and resolving the longstanding $J/\psi$ polarization puzzle.

### v2

@@ IV Summary
Within the NRQCD factorization framework, we systematically investigate $\gamma\gamma\to J/\psi+\gamma$ production. Our analysis employs four representative LDME sets that consistently describe $J/\psi$ hadroproduction data, with QCD corrections applied specifically to the direct-photon channel. The results demonstrate absolute dominance of the direct process, exceeding single- and double-resolved contributions by over two orders of magnitude; at CEPC’s design luminosity, this yields $\mathcal{O}(10^{3})$ reconstructed $J/\psi$ events annually. NLO corrections reduce the CS cross section by $\sim$50% while CO states show moderate contributions. For polarization, the CS model predicts pronounced transverse polarization. In sharp contrast, the NRQCD predictions exhibit strong LDME dependence: while two sets yield results consistent with the CSM, another set generates near-unpolarized to moderately longitudinal behavior, and the remaining one produces near-unpolarized to slightly transverse polarization. This spread highlights a significant theoretical divergence. Furthermore, the polarization predictions depend critically on the ratio $\frac{\langle\mathcal{O}^{J/\psi}(^{3}P_{J}^{[8]})\rangle}{\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[1]})\rangle}$, but not on the specific values of $\langle\mathcal{O}^{J/\psi}(^{1}S_{0}^{[8]})\rangle$ and $\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[8]})\rangle$. Thus, $\gamma\gamma\to J/\psi+\gamma+X$ provides a critical probe for testing LDME universality and resolving the longstanding $J/\psi$ polarization puzzle.